OpenStudy (anonymous):

Hello :) I need help.... 2x+5=6x-15 -5 -5 2x=6x-20 .......What do I do now???

jimthompson5910 (jim_thompson5910):

2x=6x-20 2x-6x=-20 -4x=-20 x=-20/(-4) x=5

OpenStudy (anonymous):

- 6x from both sides. or to keep it clean, +20 and -2x to both sides.

OpenStudy (anonymous):

How did you get 4x?

jimthompson5910 (jim_thompson5910):

subtract 6x from 4x to get 2x-6x=(2-6)x=-4x

OpenStudy (anonymous):

Don't you add 6 to get rid of it?

jimthompson5910 (jim_thompson5910):

I subtracted 6x from both sides to move 6x from the right side to the left side.

OpenStudy (anonymous):

The 6x is positive. You need to subtract 6x so it = 0

OpenStudy (anonymous):

2x=6x-20 2x-6x=-20 -4x=-20 x=-20/-4 x=5

OpenStudy (anonymous):

2x=6x-20 2x-6x=-20 -4x=-20 x=-20/-4 x=5

jimthompson5910 (jim_thompson5910):

In more detail 2x=6x-20 2x-6x=6x-20-6x 2x-6x=(6x-6x)-20 2x-6x=0x-20 2x-6x=-20

OpenStudy (anonymous):

How i'd do it... (I like to keep pellet positive) 2x=6x-20 0 = 4x -20 ... I -2x from both sides 20 = 4x ... I +20 to both sides. 20/4 = x ... I divided both sides by 4 x = 5

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{U}}\color{orange}{\text{N}}\color{#9c9a2e}{\text{I}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{R}}\color{purple}{\text{N}}\color{red}{\text{\S}}\color{orange}{\text{ }}\color{#9c9a2e}{\text{a}}\color{green}{\text{n}}\color{blue}{\text{d}}\color{purple}{\text{ }}\color{purple}{\text{R}}\color{red}{\text{A}}\color{orange}{\text{I}}\color{#9c9a2e}{\text{N}}\color{green}{\text{B}}\color{blue}{\text{O}}\color{purple}{\text{W}}\color{purple}{\text{\S}}\color{red}{\text{!}}\color{orange}{\text{ }}\color{#9c9a2e}{\text{}}\end{array}\]

OpenStudy (anonymous):

OMG! Hi Saifooo!!!! Thank you for the rainbowwww!!!!!!! xD

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

Heyy!

OpenStudy (anonymous):

Welcome back!!! :,)

OpenStudy (anonymous):

Lol...that makes it more fun...Math is wizard... Me + Math = Love/hate relationship